Two spherical objects have masses of 8000kg and 10.0 kg. Their centers are separated by a distance of 2.0m. Find the gravitational attraction between them.


1
Expert's answer
2021-04-19T17:21:51-0400

According to Newton's law of universal gravitation, gravitational force of interaction between two objects of mass "m_1, m_2", separated by the distance "r" from each other is "F = G \frac{m_1 m_2}{r^2}" ("G = 6.67 \cdot 10^{-11} m^3 kg^{-1} s^{-2}" is the gravitational constant).

Hence, "F = \frac{ 6.67 \cdot 10^{-11} m^3 kg^{-1} s^{-2} \cdot 8000 kg \cdot 10 kg }{(2 m)^2} \approx 1.3 \cdot 10^{-6} N"